How many gallons of an 80% alcohol solution must be mixed with 50 gallons of a 16% solution to obtain a solution that is 70% alcohol
Answer by stanbon(75887)

How many gallons of an 80% alcohol solution must be mixed with 50 gallons of a 16% solution to obtain a solution that is 70% alcohol
--------
Equation:
alcohol + alcohol = alcohol
0.80x + 0.16*50 = 0.70(x+50)
80x + 16*50 = 70x + 70*50
10x = 54*50
x = 5*54
x = 270 gallons (amt. of 80% solution needed)
==================================================
Cheers,
Stan H.
===============